### v1

# Preference-Guided Diffusion for Multi-Objective Offline Optimization

Abstract
Offline multi-objective optimization aims to identify Pareto-optimal solutions given a dataset of designs and their objective values. In this work, we propose a preference-guided diffusion model that generates Pareto-optimal designs by leveraging a classifier-based guidance mechanism. Our guidance classifier is a preference model trained to predict the probability that one design dominates another, directing the diffusion model toward optimal regions of the design space. Crucially, this preference model generalizes beyond the training distribution, enabling the discovery of Pareto-optimal solutions outside the observed dataset. We introduce a novel diversity-aware preference guidance, augmenting Pareto dominance preference with diversity criteria. This ensures that generated solutions are optimal and well-distributed across the objective space, a capability absent in prior generative methods for offline multi-objective optimization. We evaluate our approach on various continuous offline multi-objective optimization tasks and find that it consistently outperforms other inverse/generative approaches while remaining competitive with forward/surrogate-based optimization methods. Our results highlight the effectiveness of classifier-guided diffusion models in generating diverse and high-quality solutions that approximate the Pareto front well.
footnotetext: Author email: {yashas.annadani, stefan.bauer}@helmholtz-munich.de, {syrineb, ermon, bengelhardt}@stanford.edu
1 Introduction
Several design problems in science and engineering require optimizing a black-box, expensive-to-evaluate function. For example, in antibiotic drug discovery, the goal is to identify novel molecules with high antibacterial activity [31]. This can be formulated as a single-objective optimization problem. However, in practice, most real-world design challenges involve balancing multiple conflicting objectives. For instance, in drug discovery, besides maximizing antibacterial activity, we also aim to minimize toxicity and production costs [30]. This constitutes a multi-objective experimental design problem.
Prior work in both single and multi-objective optimization (MOO) has largely focused on adaptive experimental design using online methods such as Bayesian optimization [25]. These approaches rely on training surrogate models for each objective function and designing acquisition functions that are typically optimized via gradient-based techniques [3] or evolutionary algorithms to determine the next candidate for evaluation. This process is iteratively repeated to optimize the objectives. However, in many real-world applications, sequential evaluations—where inputs are tested one at a time or in small batches—are impractical. In some cases, we have only a single opportunity to evaluate the function, and, we must allocate the entire evaluation budget efficiently.
For example, in drug design, scientists cannot test molecules one by one in wet lab experiments due to the high cost, slow turnaround, and the inherently parallelizable nature of the process [30]. Instead, it is common to evaluate all candidate molecules in a single batch. This setting is referred to as offline black-box optimization [33]. While recent work has explored offline optimization in the single-objective setting [38, 18], where extensive prior data is leveraged to model the objective function and identify potential optima, the MO case remains relatively underexplored.
Offline black-box optimization presents distinct challenges compared to traditional online optimization. Since the algorithm cannot iteratively refine the learned model using newly acquired data, it must effectively leverage the available dataset to generalize beyond observed data points. This is particularly challenging because the true optima are often expected to lie outside the existing dataset, requiring robust extrapolation. Additionally, the goal is not merely to identify data points with high function values but to find solutions that satisfy a well-defined notion of optimality, such as Pareto optimality, in multi-objective settings.
Prior work in single-objective offline optimization can generally be categorized into two main approaches. Forward approaches attempt to mimic strategies used in online optimization while leveraging offline data [33]. These methods train a surrogate model of the objective function and optimize it using gradient-based techniques to propose a set of promising inputs for evaluation. While effective when the search space is well-defined, forward methods rely on having a known set of candidate inputs to evaluate. In contrast, inverse approaches use generative models to learn an inverse mapping from function values to inputs, enabling the generation of new candidates with potentially high objective values [20, 5, 13, 19]. This distinction is critical in many real-world science and engineering problems where the optimal inputs are not known in advance. For example, in chemistry, if the goal is to evaluate a known molecule, surrogate models are effective in predicting its properties. However, if the goal is to discover entirely new molecules with desired properties, inverse methods are essential, as they directly generate novel candidates rather than selecting from a predefined space. Some generative modeling approaches also draw inspiration from online methods by constructing synthetic optimization trajectories from offline data, aiming to generate new optimal points by extrapolating from the learned trajectories [18].
Multi-objective offline optimization introduces additional challenges beyond those encountered in the single-objective setting. In single-objective optimization, the goal is simply to maximize (or minimize) a function value. However, in the multi-objective case, optimality is defined in terms of Pareto optimality, which seeks the best trade-offs among competing objectives. A solution is considered Pareto optimal in the multi-objective setting when no other solution in the search space dominates the Pareto optimal solution across all objectives, meaning that improving one objective would necessarily degrade another [4]. The Pareto front includes the objective values for all Pareto-optimal points (the Pareto set).
Beyond identifying Pareto-optimal solutions, another critical challenge is ensuring diversity on the Pareto front. A well-structured Pareto front should provide solutions spread across different regions of the objective space, representing a broad range of Pareto-optimal designs. If all high-performing solutions are concentrated in a narrow region of the Pareto front, the optimization process fails to capture the full spectrum of desirable trade-offs, limiting its usefulness in real-world decision-making. This issue of diversity is well recognized even in online multi-objective optimization [1], where iterative refinement can be used to adjust the design choices. In the offline setting, however, ensuring both optimality and diversity becomes even more challenging, as the algorithm must infer these solutions solely from pre-existing data without the ability to iteratively refine its choices.
Xue et al. [37] has recently explored benchmarking offline multi-objective optimization (MOO) by building offline datasets for a variety of MOO benchmarks and proposing several potential algorithms that can provide simple solutions. Their work extends some offline single-objective optimization (SOO) approaches to the MOO setting, such as fitting surrogate models to the offline data and optimizing over the surrogates using evolutionary algorithms. However, this benchmarking paper did not tackle or suggest any approach on how to use generative models for offline MOO.
In this work, we propose a novel algorithm that leverages diffusion models for offline MOO. Our approach formulates the problem as an inverse problem, using the generative capabilities of diffusion models to produce diverse and high-quality candidate solutions that extend beyond the regions of space covered by the training data. Diffusion models have demonstrated a strong ability to generate novel samples distinct from the training data [11], making them well-suited for this task.
To incorporate Pareto optimality into the generation process, we introduce a preference-based classifier that guides the diffusion model towards Pareto-optimal solutions. This classifier is trained to compare two candidate designs and determine which one is more likely to dominate the other in the objective space. By integrating this preference-based guidance into the sampling process, we steer the generative model toward regions containing Pareto-optimal solutions.
Additionally, we address the diversity challenge in the Pareto front generation. Rather than merely identifying high-function-value designs, we train the classifier to favor solutions that are not only optimal but also well-distributed across the objective space. To achieve this, we incorporate a crowding distance-based criterion into the preference pairs used for training. Solutions with higher crowding distance, indicating greater separation from neighboring points, are considered more diverse and are prioritized in the optimization process. This ensures that the resulting Pareto front is not overly concentrated in a limited region but instead captures a wide range of trade-offs between objectives.
2 Background
2.1 Offline Multi-Objective Optimization
Multi-objective optimization (MOO) seeks to find the design ${\bm{x}}\in{\mathcal{X}}^{d}$ that minimizes (or maximizes) a set of $m$ different objectives:
$\min_{{\bm{x}}\in{\mathcal{X}}^{d}}{\bm{f}}({\bm{x}})\coloneqq\{f_{1}({\bm{x}}),\dots,f_{m}({\bm{x}})\},$ (1)
where $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$ is an unknown and expensive to evaluate objective. For most practical problems, the objectives are not simultaneously optimizable by a single design. Hence, the goal instead is to find the set of designs that are Pareto optimal.
Definition 2.1 (Pareto Dominance).
A design ${\bm{x}}$ Pareto dominates another design ${\bm{z}}\in{\mathcal{X}}^{d}$ (denoted by ${\bm{x}}\prec{\bm{z}}$) if $f_{i}({\bm{x}})\leq f_{i}({\bm{z}})\,\,\forall i$ and $\exists j:f_{j}({\bm{x}})<f_{j}({\bm{z}})$.
Definition 2.2 (Pareto Optimality and Pareto Front).
A design ${\bm{x}}^{*}$ is Pareto optimal if $\nexists\,{\bm{x}}\in{\mathcal{X}}^{d}$ such that ${\bm{x}}\prec{\bm{x}}^{*}$. The set of all Pareto optimal designs is called a Pareto set. Correspondingly, the objective values of the Pareto set $\{{\bm{f}}({\bm{x}}^{*})\mid{\bm{x}}^{*}\}$ is called the Pareto front (PF).
The Pareto front provides an optimal set of trade-offs that can be achieved from the objectives when they are not simultaneously optimizable.
Sequential methods that collect data by selecting designs and evaluating their function values are the most common approach for MOO, making use of surrogate models with uncertainty quantification to learn the unknown objectives. However, for many practical problems, these sequential methods are not feasible due to prohibitive cost or time constraints (or both). Instead of iteratively allocating an evaluation budget to refine the design choices, offline optimization uses the entire budget in a single round of function evaluations. In offline optimization, we have access to a dataset of $N$ non-optimal design-objective values pairs ${\mathcal{D}}:=\{({\bm{x}}^{(i)},{\bm{f}}({\bm{x}}^{(i)})\}_{i=1}^{N}$. The goal of offline MOO is to find the Pareto-optimal set by relying only on the existing dataset ${\mathcal{D}}$.
2.2 Diffusion Models
Diffusion models [26, 14, 27] are a class of generative models defined by a Markov chain that sequentially adds noise to data samples and then learning to denoise it by reversing the Markov chain. In this work, we follow the Denoising Diffusion Probabilistic Models (DDPM) [14] approach, which we summarize here. Given a sample ${\bm{x}}_{0}\sim q({\bm{x}})$, a time-dependent forward noising process is defined as:
$q({\bm{x}}_{t}\mid{\bm{x}}_{t-1})\coloneqq{\mathcal{N}}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},\beta_{t}{\mathbb{I}}_{d}),$ (2)
where $\beta_{t}$ is the variance of the noising schedule at timestep $t$ such that $\beta_{1}<\beta_{2}<\dots<\beta_{T}$, and $T$ is the total number of timesteps. Let $\alpha_{t}=1-\beta_{t}$ and $\tilde{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$; then the noised sample ${\bm{x}}_{t}$ can be obtained in closed form:
${\bm{x}}_{t}=\sqrt{\tilde{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\tilde{\alpha}_{t}}\epsilon_{t},\,\,\,\,\,\,\epsilon_{t}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d}).$ (3)
The reverse process conditional $q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})$ is not tractable. Therefore, a denoising model defined as $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})\coloneqq{\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}({\bm{x}}_{t}),\Sigma_{\theta}({\bm{x}}_{t}))$ is learned by optimizing parameters $\theta$. Instead of parameterizing the reverse process to estimate $\mu_{\theta}({\bm{x}}_{t})$, it is common to reparameterize this reverse process to predict the noise that was added to produce ${\bm{x}}_{t}$ (equation 3). If $\epsilon_{\theta}({\bm{x}}_{t},t)$ is the denoising model to predict the added noise, reparameterization yields the mean $\mu_{\theta}({\bm{x}}_{t})$ as:
$\mu_{\theta}({\bm{x}}_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left({\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}({\bm{x}}_{t},t)\right).$ (4)
The reverse process variance is set to be the same as the forward process variance at time $t$, i.e., $\Sigma_{\theta}({\bm{x}}_{t})\coloneqq\beta_{t}{\mathbb{I}}_{d}$. The denoising model can be trained with mean squared error loss:
$\ell(\theta)=\operatorname*{\mathbb{E}}_{t,\epsilon_{t},{\bm{x}}_{0}}\left[\left\lVert\epsilon_{t}-\epsilon_{\theta}({\bm{x}}_{t},t)\right\rVert_{2}^{2}\right].$ (5)
A new sample can be generated by first sampling $\tilde{\bm{x}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$ and then autoregressively sampling from ${\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}(\tilde{{\bm{x}}_{t}}),\beta_{t}{\mathbb{I}}_{d})$ to get $\tilde{\bm{x}}_{0}$.
2.3 Classifier Guidance
If label ${\bm{y}}$ corresponding to each sample ${\bm{x}}_{0}$ is available, then classifier guidance allows one to generate new samples from a trained diffusion model specific to a desired label. To do this, classifier guidance [11] trains an additional time-dependent classifier of the input $p_{\phi}({\bm{y}}\mid{\bm{x}}_{t},t)$. Along with the trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$, a new conditional sample can be generated by first sampling $\tilde{\bm{x}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$ and then, from $t=T$ to $t=1$, autoregressively sampling:
$\tilde{\bm{x}}_{t-1}\mid{\bm{y}},\tilde{\bm{x}}_{t}\sim{\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}(\tilde{\bm{x}}_{t})+w\beta_{t}\nabla_{\tilde{\bm{x}}_{t}}\log p_{\phi}({\bm{y}}\mid\tilde{\bm{x}}_{t},t),\beta_{t}{\mathbb{I}}_{d}),$ (6)
where $w$ is the guidance strength. Classifier guidance has been successfully used in various domains including computer vision and audio to generate samples from a certain class or label. However, their utility is less explored in black-box optimization. In this work, we use classifier guidance to generate samples that approximate the optimal Pareto sets.
3 Related Work
Online Multi-Objective Black-Box Optimization
Existing work in adaptive experimental design has addressed the multi-objective optimization problem primarily in an online fashion, where solutions are refined iteratively based on newly acquired data [1, 4]. While research on multi-objective optimization is less extensive than in the single-objective setting, several approaches have successfully tackled the problem sequentially.
One of the most prominent methods is Bayesian optimization (BO), which typically consists of three main components: a surrogate model, an acquisition function, and an optimizer. The surrogate model approximates the objective functions based on previously observed data, while the acquisition function, built using the surrogate model, is optimized to determine the next input to evaluate [25]. Most BO methods rely on Gaussian processes [36], particularly in data-scarce scenarios where no prior data is available. Some approaches simplify the multi-objective problem by reducing it to a single-objective formulation using scalarization techniques, such as linear scalarization or Chebyshev scalarization [16]. More advanced methods have leveraged expected hypervolume improvement (EHVI) [12] or information gain-based acquisition functions [4] to better navigate the trade-offs between competing objectives.
Recent advances in BO have also introduced batch selection strategies improving efficiency in parallelizable experiments [9]. However, while most of these approaches focus on identifying Pareto-optimal solutions, a few methods including Konakovic Lukovic et al. [17], Ahmadianshalchi et al. [1] explicitly address the diversity of the Pareto front. Instead, the majority of existing techniques prioritize Pareto dominance while neglecting the distribution of solutions across the objective space.
Beyond small models, some approaches have explored the use of neural networks to extend online MOO techniques by leveraging existing data. These methods often incorporate generative models, such as variational autoencoders (VAEs), combined with Gaussian processes over the latent space of the VAE [28]. However, these techniques inherit VAE’s well-known challenges, including posterior collapse and unidentifiability, which can limit their effectiveness in practical optimization settings.
Offline Single Objective Black-Box Optimization
Several approaches have been proposed to tackle single-objective offline optimization. Forward approaches, such as Trabucco et al. [33], Yu et al. [38], train a surrogate model to approximate the objective function and then optimize this model to identify a set of inputs with potentially optimal performance. Other methods [20, 5, 13] adopt generative adversarial networks (GANs) with conditional generation to learn an inverse mapping from the function space to the input space. However, these approaches inherit the well-known challenges of GANs, including unstable training dynamics and mode collapse, which can hinder their reliability in practical applications.
To address the limitations of standard forward and inverse approaches, Chemingui et al. [6], Krishnamoorthy et al. [18] proposed methods that mimic online optimization by constructing synthetic optimization trajectories from offline data. This approach trains a sequence-based model to generate new candidate solutions by extrapolating from learned optimization paths.
More recently, Krishnamoorthy et al. [19] introduced a diffusion-based optimization framework, modifying the diffusion model’s loss function to incorporate weighted importance terms. This adjustment encourages the model to prioritize inputs with higher function values while penalizing those with lower values, effectively guiding the generative process toward promising regions of the input space.
Additionally, Trabucco et al. [34] proposed a benchmarking framework to standardize and facilitate the evaluation of offline optimization algorithms, providing a consistent platform for comparing different approaches.
Offline Multi-Objective Black-Box Optimization
There is very limited prior work on offline multi-objective optimization (MOO). Xue et al. [37] recently explored this area by introducing a benchmarking framework, where they constructed offline datasets for various MOO benchmark problems and proposed several baseline algorithms. Their approach primarily involved extending forward techniques from offline single-objective optimization (SOO) to the multi-objective setting, such as fitting surrogate models to offline data and optimizing over these surrogates using evolutionary algorithms. However, this work did not explore or propose methods for leveraging generative models for inverse problems in offline MOO.
A concurrent study by Yuan et al. [40] introduces a flow-based generative model for offline MOO. Their approach scalarizes the objective functions by incorporating a weighting scheme directly into the loss function of the generative model. While this method provides an inverse method perspective, it does not address the diversity challenge in Pareto front generation, which is a critical aspect of multi-objective optimization.
4 Method: Preference-Guided Diffusion for Offline Multi-Objective Optimization
We present a new effective approach for offline MOO by using classifier guidance to generate samples from Pareto optimal sets with a diffusion model trained on offline data. Our approach does not require training individual surrogate models for each objective. It relies on an inverse strategy while ensuring the ability to generate diverse samples from the Pareto optimal set. We refer to our method as Preference-Guided Diffusion for Multi-objective Offline optimization (PGD-MOO)
Let ${\bm{x}}\in{\mathcal{X}}^{d}\subseteq{\mathbb{R}}^{d}$ be any $d$-dimensional design with corresponding objective values $y_{i}=f_{i}({\bm{x}})$ defined by unknown and expensive to evaluate functions $f_{i}:{\mathcal{X}}^{d}\mapsto{\mathbb{R}}$. Let ${\bm{y}}=[y_{1},\dots,y_{m}]^{T}$ be the vector of objective values for an $m$-objective problem. In offline MOO, we have access to a dataset ${\mathcal{D}}:=\{{\bm{x}}^{(i)},{\bm{y}}^{(i)}\}_{i=1}^{N}$ of $N$ previously evaluated design-objective pairs. Given ${\mathcal{D}}$, the goal is to generate designs ${\bm{x}}^{*}$ from the unknown optimal Pareto set.
While diffusion models capture the distribution over data $p({\bm{x}})$, in offline MOO, we are often interested in samples that lie outside the training data, closer to the Pareto front. This motivates the use of classifier guidance. Directly using classifier guidance in diffusion models usually involves training surrogate models for each objective, which often requires scalarization and hence can be suboptimal. We propose to use preference-based guidance to capture Pareto dominance relations between data points.
Figure 1: Generalization of the preference model on regions unseen in the training data on the ZDT2 task [45]. The preference model gives good prediction of Pareto dominance between reference design (in black) with other designs. The figure on the right is a zoomed-in version of the left, excluding the training data (in blue).
4.1 Preference Guided Diffusion
In this work, we explore an alternate guidance strategy that does not involve training surrogate models for every objective. Instead, we train a preference model that predicts whether a design Pareto dominates (Definition 2.1) another design. Given two designs ${\bm{x}}$ and $\hat{\bm{x}}$, we train a (time-conditioned) binary classifier that predicts $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$. We parameterize this distribution with a multi-layer perceptron (MLP) that takes in two inputs (designs) of size $2\times d$ and outputs the logit of the Bernoulli distribution predicting whether the first input Pareto dominates the second input.
Training the Preference Classifier.
To train the preference classifier, we first sort the points in the training data by their Pareto dominance. The data is then divided into multiple fronts in increasing order of dominance, with points in the same front considered equally dominant. Next, We select $({\bm{x}},\hat{\bm{x}})$ pairs randomly from the dataset. If one design strictly dominates the other, it is labeled as preferred. Otherwise, if there is no strict dominance between either of the selected designs $({\bm{x}},\hat{\bm{x}})$, we assign the label ${\bm{x}}\prec\hat{\bm{x}}$ if ${\bm{x}}$ has more diversity contribution than $\hat{\bm{x}}$ wrt the other designs in the dataset that belong to the same Pareto front (and vice versa). In this work, we calculate the diversity contribution of each point by computing the crowding distance [10] of a selected design w.r.t all other points that belong to the same front in the dataset. Crowding distance for any point ${\bm{x}}$ is computed as follows:
$d_{\text{CD}}({\bm{x}})=\sum_{i=1}^{m}\frac{y_{i}^{+}-y_{i}^{-}}{y_{i}^{{\mathcal{D}}(\text{max})}-y_{i}^{{\mathcal{D}}(\text{min})}},$ (7)
where $y_{i}^{+}$ and $y_{i}^{-}$ are the $i$th objective values of neighboring designs of ${\bm{x}}$ in the corresponding front sorted according to $i$th objective value. $f_{i}^{{\mathcal{D}}(\text{max})}$ and $f_{i}^{{\mathcal{D}}(\text{min})}$ are maximum and minimum values of objective $i$ in the entire dataset. Crowding distance has been used as a secondary selection criterion in evolutionary algorithms like NSGA-2 [10] to maintain diversity of solutions. Using crowding distance to create a binary label encourages the preference model to not only guide the diffusion model towards more Pareto-dominant regions but also ensure that the designs that make up the Pareto front are diverse.
For the denoising model, we train an unconditional diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ (§ 2.2) on the designs ${\bm{x}}$ in the dataset, similar to DDPM [14].
Algorithm 1 Sampling from Preference Guided Diffusion
0:  Trained $\epsilon_{\theta}({\bm{x}}_{t},t)$, preference model $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$, guidance weight $w$ and the most dominant design in the dataset ${\bm{x}}^{{\mathcal{D}}(\text{best})}$
1:  $r\leftarrow{\bm{x}}^{{\mathcal{D}}(\text{best})}$
2:  $\tilde{{\bm{x}}}_{T}\sim{\mathcal{N}}(0,{\mathbb{I}}_{d})$
3:  for t = $T$ to $1$ do
4:     $\mu_{\theta}(\tilde{\bm{x}}_{t})=\frac{1}{\sqrt{\alpha_{t}}}\left(\tilde{\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}(\tilde{\bm{x}}_{t},t)\right)$
5:     Compute preference score $s_{p}=\nabla_{\tilde{\bm{x}}_{t}}p_{\phi}(\tilde{\bm{x}}_{t}\prec r\mid\tilde{\bm{x}}_{t},r,t)$
6:     Sample $\tilde{\bm{x}}_{t-1}\sim{\mathcal{N}}({\bm{x}}_{t-1};\mu_{\theta}(\tilde{\bm{x}}_{t})+w\beta_{t}s_{p},\beta_{t}{\mathbb{I}}_{d})$
7:     $r\leftarrow\tilde{\bm{x}}_{t}$
8:  end for
9:  return  $\tilde{\bm{x}}_{0}$
Sampling Designs.
With a trained denoising model $\epsilon_{\theta}({\bm{x}}_{t},t)$ and preference model $p_{\phi}({\bm{x}}\prec\hat{\bm{x}}\mid{\bm{x}},\hat{\bm{x}},t)$, we sample a new design by using classifier guidance (§ 2.3). We input both the denoised variable at the current timestep $\tilde{\bm{x}}_{t}$ as well as from the previous realization of denoising, i.e., $\tilde{\bm{x}}_{t+1}$ for the preference model to estimate $\nabla_{\tilde{\bm{x}}_{t}}p_{\phi}(\tilde{\bm{x}}_{t}\prec\tilde{\bm{x}}_{t+1}\mid\tilde{\bm{x}}_{t},\tilde{\bm{x}}_{t+1},t)$. The sampling procedure is summarized in Algorithm 1. Intuitively, a preference model that generalizes well beyond its training data should guide the denoising process such that, at each step of denoising, the resulting sample $\tilde{\bm{x}}_{t}$ is in a more Pareto-dominant region. With enough timesteps, the resulting denoised sample $\tilde{\bm{x}}_{0}$ will be close to the Pareto front. This approach does not require training surrogate models, thus providing a simple alternative approach to offline MOO. We find that the preference model generalizes well outside of the training data (see Fig. 1), therefore providing guidance to the diffusion model to generate designs outside of the training data close to the Pareto front, while maintaining diversity in the samples.
5 Experiments
We perform several experiments on standard benchmarks for offline MOO. Through these experiments, we would like to understand how close the generated samples are to the Pareto front, as well as the diversity of the solutions.
Benchmark Tasks
Our evaluation closely follows the benchmarking effort provided in prior work  [37]. We evaluate our approach on two sets of tasks - synthetic and real-world applications-based RE engineering suite [32].
  1. 1.
    Synthetic tasks consist of several subtasks wherein the objective functions are hand-designed with known ground truth Pareto-fronts. These subtasks have been widely used in MOO problems to study the performance of the algorithm. These subtasks consist of 2-3 objectives with $d$ ranging from 10 to 30. We use the same dataset as prior work [37], which consists of 60,000 offline data points.
  2. 2.
    RE engineering suite of problems are set of tasks that are based on real-world applications in engineering, for instance ,rocket injector design and disc brake design. $d$ ranges from 3-7 variables and the number of objectives $m$ varies from 2 to 6. We use the same dataset as in prior work [37], which consists of 60,000 offline data points.
Baselines
We primarily compare our approach with two categories of baselines:
  1. 1.
    We compare with ParetoFlow [40], a classifier guided generative model based on flow-matching [22]. Classifiers for guidance are trained surrogate models for each objective, followed by scalarization. It is important to note that ParetoFlow is our primary baseline since we are targeting inverse methods. In an inverse problem setting, the forward approaches introduced next are not applicable.
  2. 2.
    Forward approaches using evolutionary algorithms: As suggested in prior work [37], a standard approach to offline MOO is to train a surrogate model for each objective and then use evolutionary algorithms such as NSGA-2 [10] to search over the design space. Although there are various ways to learn surrogate models, we compare with deep neural network (DNN)-based approaches, which are shown to perform best according to benchmarks [37]. The DNN approaches we compare with are: i) A Multi-Head Model: Uses multi-task learning [42] to train a joint surrogate for all objectives. Training techniques for this approach such as GradNorm [8] and PcGrad [39] are also compared. ii) Multiple Models: Maintain $m$ independent surrogate models, each making use of a single optimization technique, including COMs [33], ROMA [38], IOM [24], ICT [41], and Tri-mentoring [7].
Evaluation Metrics
For each algorithm, we evaluate the convergence of solutions using the hypervolume metric [44], a standard metric in MOO for measuring the closeness of the proposed designs to the Pareto front. Hypervolume measures the volume of the objective space between a reference point and the objective vectors of the solution set, and does not require access to the true Pareto front.
In addition to the hypervolume, we also measure the diversity of the obtained solutions using the $\Delta$-spread metric [10, 43]. The $\Delta$-spread measures the extent of the spread achieved in a computed Pareto front approximation [2]. It is important to consider the diversity of the obtained solutions, especially in the case of MOO wherein there is no single “best” design, but rather an entire set of solutions based on the Pareto front. In addition, in the case of offline optimization, the acquisition is single-shot. Therefore, solutions that are diverse and hence provide more coverage over the objective space are preferable. In this work, we provide the first effort to evaluate and benchmark the diversity of solutions obtained by different approaches in offline MOO. We evaluate all methods on five random seeds, and we compute the metrics using a budget of 256 designs.
Figure 2: Plot of the samples from diffusion model (in green) on the ZDT2 task [45] at different timesteps of denoising. Top row shows results from our preference-guided diffusion, while the bottom row shows results from a purely unconditional model. Convergence of samples close to the Pareto front (in red) outside of the training data (blue) highlights the importance of preference guidance.
Table 1: Hypervolume results of DTLZ subtasks (part of the synthetic task). Each method is run for five random seeds and evaluated on 256 designs.
Method DTLZ1 DTLZ2 DTLZ3 DTLZ4 DTLZ5 DTLZ6 DTLZ7
D (best) 10.60 9.91 10.00 10.76 9.35 8.88 8.56
MultiHead 10.51 $\pm$ 0.23 9.03 $\pm$ 0.56 10.48 $\pm$ 0.23 6.73 $\pm$ 1.4 8.41 $\pm$ 0.15 8.72 $\pm$ 1.07 10.66 $\pm$ 0.09
MultiHead - PcGrad 10.64 $\pm$ 0.01 9.64 $\pm$ 0.33 10.55 $\pm$ 0.12 9.95 $\pm$ 1.93 9.02 $\pm$ 0.24 9.90 $\pm$ 0.25 10.61 $\pm$ 0.03
MultiHead - GradNorm 10.64 $\pm$ .01 8.86 $\pm$ 1.27 10.26 $\pm$ 0.28 7.45 $\pm$ 0.75 7.87 $\pm$ 1.06 8.16 $\pm$ 2.21 10.31 $\pm$ 0.22
MultipleModels 10.64 $\pm$ 0.01 9.03 $\pm$ 0.80 10.58 $\pm$ 0.03 7.66 $\pm$ 1.3 7.65 $\pm$ 1.39 9.58 $\pm$ 0.31 10.61 $\pm$ 0.16
MultipleModels - COM 10.64 $\pm$ 0.01 8.99 $\pm$ 0.97 10.27 $\pm$ 0.37 9.72 $\pm$ 0.39 9.44 $\pm$ 0.41 9.37 $\pm$ 0.35 10.09 $\pm$ 0.36
MultipleModels - IOM 10.64 $\pm$ 0.01 10.10 $\pm$ 0.27 10.24 $\pm$ 0.13 10.03 $\pm$ 0.53 9.77 $\pm$ 0.18 9.30 $\pm$ 0.31 10.60 $\pm$ 0.05
MultipleModels - ICT 10.64 $\pm$ 0.01 8.68 $\pm$ 0.88 10.25 $\pm$ 0.42 10.33 $\pm$ 0.24 9.25 $\pm$ 0.28 9.10 $\pm$ 1.16 10.29 $\pm$ 0.05
MultipleModels - RoMA 10.64 $\pm$ 0.01 10.04 $\pm$ 0.05 10.61 $\pm$ 0.03 9.25 $\pm$ 0.11 8.71 $\pm$ 0.47 9.84 $\pm$ 0.25 10.53 $\pm$ 0.04
MultipleModels - TriMentoring 10.64 $\pm$ 0.01 9.39 $\pm$ 0.35 10.48 $\pm$ 0.12 10.21 $\pm$ 0.06 7.69 $\pm$ 1.03 9.00 $\pm$ 0.48 10.12 $\pm$ 0.09
ParetoFlow 10.60 $\pm$ 0.02 10.13 $\pm$ 0.16 10.41 $\pm$ 0.09 10.29 $\pm$ 0.17 9.65 $\pm$ 0.23 9.25 $\pm$ 0.43 8.94 $\pm$ 0.18
PGD-MOO + Data Pruning (Ours) 10.64 $\pm$ 0.01 10.55 $\pm$ 0.01 10.63 $\pm$ 0.01 10.63 $\pm$ 0.01 10.07 $\pm$ 0.02 10.15 $\pm$ 0.03 9.57 $\pm$ 0.07
PGD-MOO (Ours) 10.65 $\pm$ 0.01 10.55 $\pm$ 0.01 10.63 $\pm$ 0.01 10.64 $\pm$ 0.01 10.06 $\pm$ 0.02 10.14 $\pm$ 0.01 9.70 $\pm$ 0.18
Table 2: Hypervolume results of ZDT subtasks (part of the synthetic task) along with average rank of each method on the entire synthetic set of tasks. Each method is run for five random seeds and evaluated on 256 designs.
Method ZDT1 ZDT2 ZDT3 ZDT4 ZDT6 Avg. Rank
D (best) 4.17 4.67 5.15 5.45 4.61 8.43
MultiHead 4.8 $\pm$ 0.03 5.57 $\pm$ 0.07 5.58 $\pm$ 0.2 4.59 $\pm$ 0.26 4.78 $\pm$ 0.01 7.5
MultiHead - PcGrad 4.84 $\pm$ 0.01 5.55 $\pm$ 0.11 5.51 $\pm$ 0.03 3.68 $\pm$ 0.70 4.67 $\pm$ 0.1 6.08
MultiHead - GradNorm 4.63 $\pm$ 0.15 5.37 $\pm$ 0.17 5.54 $\pm$ 0.2 3.28 $\pm$ 0.9 3.81 $\pm$ 1.2 9.75
MultipleModels 4.81 $\pm$ 0.02 5.57 $\pm$ 0.07 5.48 $\pm$ 0.21 5.03 $\pm$ 0.19 4.78 $\pm$ 0.01 6.5
MultipleModels - COM 4.52 $\pm$ 0.02 4.99 $\pm$ 0.12 5.49 $\pm$ 0.07 5.10 $\pm$ 0.08 4.41 $\pm$ 0.21 7.83
MultipleModels - IOM 4.68 $\pm$ 0.12 5.45 $\pm$ 0.11 5.61 $\pm$ 0.06 4.99 $\pm$ 0.21 4.75 $\pm$ 0.01 5.58
MultipleModels - ICT 4.82 $\pm$ 0.01 5.58 $\pm$ 0.01 5.59 $\pm$ 0.06 4.63 $\pm$ 0.43 4.75 $\pm$ 0.01 6.67
MultipleModels - RoMA 4.84 $\pm$ 0.01 5.43 $\pm$ 0.35 5.89 $\pm$ 0.04 4.13 $\pm$ 0.11 1.71 $\pm$ 0.10 6.58
MultipleModels - TriMentoring 4.64 $\pm$ 0.10 5.22 $\pm$ 0.11 5.16 $\pm$ 0.04 5.12 $\pm$ 0.12 2.61 $\pm$ 0.01 8.33
ParetoFlow 4.23 $\pm$ 0.04 5.65 $\pm$ 0.11 5.29 $\pm$ 0.14 5.00 $\pm$ 0.22 4.48 $\pm$ 0.11 7.58
PGD-MOO + Data Pruning (Ours) 4.54 $\pm$ 0.08 5.21 $\pm$ 0.06 5.61 $\pm$ 0.06 5.06 $\pm$ 0.07 4.56 $\pm$ 0.14 5.08
PGD-MOO (Ours) 4.41 $\pm$ 0.08 5.33 $\pm$ 0.05 5.54 $\pm$ 0.10 5.02 $\pm$ 0.03 4.82 $\pm$ 0.01 4.42
5.1 Training Details
We parameterize the unconditional denoising model to be a multi-layer perceptron (MLP) with two 512-dimensional hidden layers, followed by a ReLU nonlinearity and layer normalization [21]. We also incorporate sinusoidal time embedding [35] for conditioning. We parameterize the preference model to be an MLP with three hidden layers, with first two hidden layers having the same number of units as the input, while the last hidden layer is having 512 units. Similar to denoising model, we also use ReLU nonlinearity followed by layer normalization and sinusoidal time embedding.
The denoising model is trained with AdamW optimizer [23] with learning rate of $45e-$ 4 for up to 200 epochs. The preference model is trained with Adam optimizer [15] with learning rate of $51e-$ 5 for up to 500 epochs. During sampling, we set the guidance weight $w$ to 10. For the preference model, we also experiment with pruning the training data to only contain the top $30\%$ of points, sorted according to their dominance. We refer to this method as PGD-MOO + Data Pruning in the results.
Table 3: Selected results of hypervolume on RE task. Results are evaluated on 256 designs with five different random seeds.
Method RE21 RE22 RE25 RE32 RE35 RE37 RE41 RE61 Avg. Rank
D(best) 4.1 4.78 4.79 10.56 10.08 5.57 18.27 97.49 11.3
MultiHead-GradNorm 4.28 $\pm$ 0.39 4.7 $\pm$ 0.44 4.52 $\pm$ 0.5 10.54 $\pm$ 0.15 9.76 $\pm$ 1.3 5.67 $\pm$ 1.41 17.06 $\pm$ 3.82 108.01 $\pm$ 1.0 11.4
MultiHead-PcGrad 4.59 $\pm$ 0.01 4.73 $\pm$ 0.36 4.78 $\pm$ 0.14 10.63 $\pm$ 0.01 10.51 $\pm$ 0.05 6.68 $\pm$ 0.06 20.66 $\pm$ 0.1 108.54 $\pm$ 0.23 6.06
MultiHead 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.74 $\pm$ 0.2 10.6 $\pm$ 0.05 10.49 $\pm$ 0.07 6.67 $\pm$ 0.05 20.62 $\pm$ 0.11 108.92 $\pm$ 0.22 5.73
MultipleModels-COM 4.38 $\pm$ 0.09 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.64 $\pm$ 0.01 10.55 $\pm$ 0.02 6.35 $\pm$ 0.1 20.37 $\pm$ 0.06 107.99 $\pm$ 0.48 7.27
MultipleModels-ICT 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.5 $\pm$ 0.01 6.73 $\pm$ 0.0 20.58 $\pm$ 0.04 108.68 $\pm$ 0.27 4.67
MultipleModels-IOM 4.58 $\pm$ 0.02 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 10.65 $\pm$ 0.0 10.57 $\pm$ 0.01 6.71 $\pm$ 0.02 20.66 $\pm$ 0.05 107.71 $\pm$ 0.5 4.6
MultipleModels-RoMA 4.57 $\pm$ 0.0 4.61 $\pm$ 0.51 4.83 $\pm$ 0.01 10.64 $\pm$ 0.0 10.53 $\pm$ 0.03 6.67 $\pm$ 0.02 20.39 $\pm$ 0.09 108.47 $\pm$ 0.28 7.67
MultipleModels-TriMentoring 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.62 $\pm$ 0.01 10.59 $\pm$ 0.0 6.73 $\pm$ 0.01 20.68 $\pm$ 0.04 108.61 $\pm$ 0.29 4.3
MultipleModels 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.63 $\pm$ 0.25 10.62 $\pm$ 0.02 10.55 $\pm$ 0.01 6.73 $\pm$ 0.03 20.77 $\pm$ 0.08 108.96 $\pm$ 0.06 3.67
ParetoFlow 4.2 $\pm$ 0.17 4.86 $\pm$ 0.01 - 10.61 $\pm$ 0.0 11.12 $\pm$ 0.02 6.55 $\pm$ 0.59 19.41 $\pm$ 0.92 107.1 $\pm$ 6.96 6.17
PGD-MOO + Data Pruning (Ours) 4.42 $\pm$ 0.04 4.83 $\pm$ 0.01 4.84 $\pm$ 0.0 10.64 $\pm$ 0.0 10.43 $\pm$ 0.04 5.99 $\pm$ 0.18 19.37 $\pm$ 0.15 103.04 $\pm$ 1.71 9.06
PGD-MOO (Ours) 4.46 $\pm$ 0.03 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.65 $\pm$ 0.0 10.32 $\pm$ 0.1 6.13 $\pm$ 0.12 19.31 $\pm$ 0.46 105.02 $\pm$ 1.14 7.67
Table 4: Average ranking of the $\Delta$-spread metric obtained by different algorithms on both synthetic and RE tasks. Detailed results are provided in Appendix A.
Method Synthetic RE
MultiHead-GradNorm 7.12 8.27
MultiHead-PcGrad 5.92 7.0
MultiHead 8.83 7.2
MultipleModels-COM 6.5 5.47
MultipleModels-ICT 6.5 5.87
MultipleModels-IOM 6.42 6.8
MultipleModels-RoMA 5.92 6.4
MultipleModels-TriMentoring 6.0 4.6
MultipleModels 9.25 7.53
ParetoFlow 10.0 9.0
PGD-MOO + Data Pruning (Ours) 2.67 4.0
PGD-MOO (Ours) 2.83 4.28
5.2 Results
Evaluation of convergence.
We provide detailed results of hypervolume for various baselines and our approach on the synthetic task (Tables 1 and 2). We find that our approach performs competitively with respect to baselines. Preference-guided diffusion performs on average better than ParetoFlow, another generative model-based approach using guidance. This shows the benefits of having a preference model as a classifier for guidance. Overall, our method performs better than other baselines, which learn surrogate models and use evolutionary algorithms in the synthetic task setting (Fig. 2). In addition, we also find that our method performs competitively in the RE engineering suite (Tables 3 and 6). In problems with higher number of objectives, we find that our approach is slightly worse compared to the baselines in terms of hypervolume. However, we note that our approach is much simpler to train in these settings, while still achieving diverse solutions (discussed further below).
Evaluation of diversity.
Average ranking in terms of performance of the $\Delta$-spread metric for all algorithms (Table 4) shows that our approach gives more diverse solutions than all the other baselines including in the RE setting. These results highlight the importance of having a diversity constraint in the training procedure for the classifier through the data selection procedure (Tables 8, 7 and 10).
Across these experiments, we find that our approach gives competitive performance in terms of hypervolume (convergence) while being better in terms of the $\Delta$-spread metric (diversity) than the baselines.
6 Conclusion
In this work, we presented a novel classifier-guided diffusion approach for offline multi-objective optimization (MOO). Our method leverages a preference model that predicts Pareto dominance between pairs of inputs, incorporating diversity considerations to ensure that designs on the same Pareto front are well-distributed. Empirical results show that our technique performs competitively in terms of convergence to the true Pareto front, while also generating a diverse set of solutions.
Limitations. A key limitation of our current approach is that it relies solely on dominance information rather than the individual function values of the objectives. Consequently, it does not allow fine-grained control over trade-offs among different objectives, which can be important if a practitioner needs to emphasize or de-emphasize specific objective values.
Future Directions. One promising extension would be to integrate additional guidance signals, such as the actual function values, enabling a more preference-based form of MOO. This would allow users to explicitly prioritize certain objectives over others or specify desired performance ranges. Another avenue for future work is combining forward (surrogate-based) and inverse (generative) approaches, where candidates proposed by the generative model are iteratively refined using surrogate models.
Acknowledgements
This work was supported by a CIFAR Catalyst grant, the Helmholtz Foundation Model Initiative and the Helmholtz Association. The authors also acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUQUEEN [29] at Jülich Supercomputing Centre (JSC). BEE was funded in part by grants from the Parker Institute for Cancer Immunology (PICI), the Chan-Zuckerberg Institute (CZI), NIH NHGRI R01 HG012967, and NIH NHGRI R01 HG013736. BEE is a CIFAR Fellow in the Multiscale Human Program. Syrine Belakaria is supported by the Stanford Data Science fellowship.
Appendix A Additional Results
Table 5: Evaluation of hypervolume with 256 sampled designs on subsets of the RE task. Results are averaged over 5 different random seeds.
Method RE21 RE22 RE23 RE24 RE25 RE31 RE32 RE33
D(best) 4.1 4.78 4.75 4.6 4.79 10.6 10.56 10.56
MultiHead-GradNorm 4.28 $\pm$ 0.39 4.7 $\pm$ 0.44 3.77 $\pm$ 1.12 3.65 $\pm$ 0.82 4.52 $\pm$ 0.5 10.6 $\pm$ 0.1 10.54 $\pm$ 0.15 10.03 $\pm$ 1.5
MultiHead-PcGrad 4.59 $\pm$ 0.01 4.73 $\pm$ 0.36 4.84 $\pm$ 0.0 4.15 $\pm$ 0.66 4.78 $\pm$ 0.14 10.64 $\pm$ 0.01 10.63 $\pm$ 0.01 10.59 $\pm$ 0.03
MultiHead 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.01 4.73 $\pm$ 0.2 4.74 $\pm$ 0.2 10.65 $\pm$ 0.0 10.6 $\pm$ 0.05 10.62 $\pm$ 0.0
MultipleModels-COM 4.38 $\pm$ 0.09 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.73 $\pm$ 0.2 4.83 $\pm$ 0.01 10.64 $\pm$ 0.01 10.64 $\pm$ 0.01 10.61 $\pm$ 0.0
MultipleModels-ICT 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.45 $\pm$ 0.02 4.83 $\pm$ 0.01 4.84 $\pm$ 0.0 10.65 $\pm$ 0.0 10.64 $\pm$ 0.0 10.62 $\pm$ 0.0
MultipleModels-IOM 4.58 $\pm$ 0.02 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 4.72 $\pm$ 0.11 4.83 $\pm$ 0.01 10.65 $\pm$ 0.0 10.65 $\pm$ 0.0 10.62 $\pm$ 0.0
MultipleModels-RoMA 4.57 $\pm$ 0.0 4.61 $\pm$ 0.51 4.83 $\pm$ 0.01 3.96 $\pm$ 1.2 4.83 $\pm$ 0.01 10.64 $\pm$ 0.01 10.64 $\pm$ 0.0 10.58 $\pm$ 0.03
MultipleModels-TriMentoring 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.65 $\pm$ 0.0 10.62 $\pm$ 0.01 10.6 $\pm$ 0.01
MultipleModels 4.6 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.83 $\pm$ 0.01 4.63 $\pm$ 0.25 10.65 $\pm$ 0.0 10.62 $\pm$ 0.02 10.62 $\pm$ 0.0
ParetoFlow 4.2 $\pm$ 0.17 4.86 $\pm$ 0.01 - - - 10.66 $\pm$ 0.12 10.61 $\pm$ 0.0 10.75 $\pm$ 0.2
PGD-MOO + Data Pruning (Ours) 4.42 $\pm$ 0.04 4.83 $\pm$ 0.01 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.57 $\pm$ 0.05 10.64 $\pm$ 0.0 10.09 $\pm$ 0.6
PGD-MOO (Ours) 4.46 $\pm$ 0.03 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 4.84 $\pm$ 0.0 10.6 $\pm$ 0.01 10.65 $\pm$ 0.0 10.51 $\pm$ 0.04
Table 6: Evaluation of hypervolume with 256 sampled designs on subsets of the RE task. Results are averaged over 5 different random seeds.
Method RE34 RE35 RE36 RE37 RE41 RE42 RE61
D(best) 9.3 10.08 7.61 5.57 18.27 14.52 97.49
MultiHead-GradNorm 8.47 $\pm$ 1.87 9.76 $\pm$ 1.3 9.67 $\pm$ 0.43 5.67 $\pm$ 1.41 17.06 $\pm$ 3.82 18.77 $\pm$ 2.99 108.01 $\pm$ 1.0
MultiHead-PcGrad 10.11 $\pm$ 0.0 10.51 $\pm$ 0.05 10.17 $\pm$ 0.08 6.68 $\pm$ 0.06 20.66 $\pm$ 0.1 22.57 $\pm$ 0.26 108.54 $\pm$ 0.23
MultiHead 10.1 $\pm$ 0.01 10.49 $\pm$ 0.07 10.23 $\pm$ 0.03 6.67 $\pm$ 0.05 20.62 $\pm$ 0.11 22.38 $\pm$ 0.35 108.92 $\pm$ 0.22
MultipleModels-COM 9.96 $\pm$ 0.09 10.55 $\pm$ 0.02 9.82 $\pm$ 0.35 6.35 $\pm$ 0.1 20.37 $\pm$ 0.06 17.44 $\pm$ 0.71 107.99 $\pm$ 0.48
MultipleModels-ICT 10.1 $\pm$ 0.0 10.5 $\pm$ 0.01 10.29 $\pm$ 0.03 6.73 $\pm$ 0.0 20.58 $\pm$ 0.04 22.27 $\pm$ 0.15 108.68 $\pm$ 0.27
MultipleModels-IOM 10.11 $\pm$ 0.01 10.57 $\pm$ 0.01 10.29 $\pm$ 0.04 6.71 $\pm$ 0.02 20.66 $\pm$ 0.05 22.43 $\pm$ 0.1 107.71 $\pm$ 0.5
MultipleModels-RoMA 9.91 $\pm$ 0.01 10.53 $\pm$ 0.03 9.72 $\pm$ 0.28 6.67 $\pm$ 0.02 20.39 $\pm$ 0.09 21.41 $\pm$ 0.37 108.47 $\pm$ 0.28
MultipleModels-TriMentoring 10.08 $\pm$ 0.02 10.59 $\pm$ 0.0 9.64 $\pm$ 1.42 6.73 $\pm$ 0.01 20.68 $\pm$ 0.04 21.6 $\pm$ 0.19 108.61 $\pm$ 0.29
MultipleModels 10.11 $\pm$ 0.0 10.55 $\pm$ 0.01 10.24 $\pm$ 0.03 6.73 $\pm$ 0.03 20.77 $\pm$ 0.08 22.59 $\pm$ 0.11 108.96 $\pm$ 0.06
ParetoFlow 11.2 $\pm$ 0.35 11.12 $\pm$ 0.02 8.42 $\pm$ 0.35 6.55 $\pm$ 0.59 19.41 $\pm$ 0.92 20.35 $\pm$ 5.31 107.1 $\pm$ 6.96
PGD-MOO + Data Pruning (Ours) 9.15 $\pm$ 0.11 10.43 $\pm$ 0.04 9.48 $\pm$ 0.33 5.99 $\pm$ 0.18 19.37 $\pm$ 0.15 17.4 $\pm$ 0.63 103.04 $\pm$ 1.71
PGD-MOO (Ours) 9.39 $\pm$ 0.16 10.32 $\pm$ 0.1 9.37 $\pm$ 0.17 6.13 $\pm$ 0.12 19.31 $\pm$ 0.46 19.01 $\pm$ 0.68 105.02 $\pm$ 1.14
Table 7: Evaluation of the $\Delta$-spread metric with 256 sampled designs on DTLZ subtask, part of the synthetic task. Results are averaged over 5 different random seeds. Lower values are better.
Method DTLZ1 DTLZ2 DTLZ3 DTLZ4 DTLZ5 DTLZ6 DTLZ7
MultiHead-GradNorm 0.61 $\pm$ 0.03 0.89 $\pm$ 0.24 0.88 $\pm$ 0.29 0.96 $\pm$ 0.14 0.75 $\pm$ 0.1 0.95 $\pm$ 0.28 1.2 $\pm$ 0.18
MultiHead-PcGrad 0.65 $\pm$ 0.04 0.73 $\pm$ 0.05 0.76 $\pm$ 0.11 1.08 $\pm$ 0.2 0.77 $\pm$ 0.06 0.87 $\pm$ 0.06 1.15 $\pm$ 0.2
MultiHead 0.86 $\pm$ 0.08 0.93 $\pm$ 0.15 0.93 $\pm$ 0.16 0.98 $\pm$ 0.12 0.92 $\pm$ 0.14 0.88 $\pm$ 0.2 0.93 $\pm$ 0.11
MultipleModels-COM 0.69 $\pm$ 0.02 0.85 $\pm$ 0.15 0.73 $\pm$ 0.06 1.09 $\pm$ 0.12 0.89 $\pm$ 0.08 1.15 $\pm$ 0.15 0.78 $\pm$ 0.03
MultipleModels-ICT 0.7 $\pm$ 0.02 0.79 $\pm$ 0.02 0.63 $\pm$ 0.1 0.92 $\pm$ 0.03 0.8 $\pm$ 0.05 0.91 $\pm$ 0.06 0.77 $\pm$ 0.05
MultipleModels-IOM 0.66 $\pm$ 0.02 0.96 $\pm$ 0.18 0.67 $\pm$ 0.04 1.23 $\pm$ 0.24 0.91 $\pm$ 0.07 1.11 $\pm$ 0.09 0.75 $\pm$ 0.03
MultipleModels-RoMA 0.62 $\pm$ 0.03 1.06 $\pm$ 0.08 0.89 $\pm$ 0.11 1.28 $\pm$ 0.08 0.93 $\pm$ 0.13 0.76 $\pm$ 0.1 0.69 $\pm$ 0.03
MultipleModels-TriMentoring 0.72 $\pm$ 0.03 0.91 $\pm$ 0.1 0.66 $\pm$ 0.09 0.95 $\pm$ 0.07 0.7 $\pm$ 0.06 0.8 $\pm$ 0.09 0.81 $\pm$ 0.05
MultipleModels 0.88 $\pm$ 0.07 1.11 $\pm$ 0.26 1.0 $\pm$ 0.22 0.93 $\pm$ 0.13 0.92 $\pm$ 0.17 1.0 $\pm$ 0.24 0.85 $\pm$ 0.12
ParetoFlow 0.82 $\pm$ 0.02 1.07 $\pm$ 0.07 0.68 $\pm$ 0.09 1.63 $\pm$ 0.15 1.04 $\pm$ 0.06 1.16 $\pm$ 0.09 0.84 $\pm$ 0.09
PGD-MOO + Data Pruning (Ours) 0.58 $\pm$ 0.04 0.52 $\pm$ 0.03 0.46 $\pm$ 0.02 1.66 $\pm$ 0.04 0.51 $\pm$ 0.02 0.63 $\pm$ 0.02 0.53 $\pm$ 0.04
PGD-MOO (Ours) 0.62 $\pm$ 0.02 0.5 $\pm$ 0.03 0.46 $\pm$ 0.02 1.6 $\pm$ 0.04 0.51 $\pm$ 0.02 0.61 $\pm$ 0.03 0.63 $\pm$ 0.02
Table 8: Evaluation of the $\Delta$-spread metric with 256 sampled designs on ZDT subtask, part of the synthetic task. Results are averaged over 5 different random seeds. Lower values are better.
Method ZDT1 ZDT2 ZDT3 ZDT4 ZDT6
MultiHead-GradNorm 0.96 $\pm$ 0.19 0.94 $\pm$ 0.16 0.93 $\pm$ 0.17 0.79 $\pm$ 0.15 0.77 $\pm$ 0.16
MultiHead-PcGrad 0.83 $\pm$ 0.1 0.98 $\pm$ 0.13 0.79 $\pm$ 0.03 0.67 $\pm$ 0.04 0.82 $\pm$ 0.11
MultiHead 1.13 $\pm$ 0.08 1.04 $\pm$ 0.05 0.83 $\pm$ 0.06 0.68 $\pm$ 0.03 1.22 $\pm$ 0.07
MultipleModels-COM 0.89 $\pm$ 0.04 0.79 $\pm$ 0.11 0.83 $\pm$ 0.05 0.64 $\pm$ 0.03 1.0 $\pm$ 0.09
MultipleModels-ICT 1.1 $\pm$ 0.02 1.01 $\pm$ 0.07 0.86 $\pm$ 0.06 0.69 $\pm$ 0.07 0.98 $\pm$ 0.06
MultipleModels-IOM 0.94 $\pm$ 0.1 0.9 $\pm$ 0.05 0.81 $\pm$ 0.07 0.73 $\pm$ 0.04 0.46 $\pm$ 0.1
MultipleModels-RoMA 0.64 $\pm$ 0.06 0.92 $\pm$ 0.1 0.79 $\pm$ 0.07 0.69 $\pm$ 0.02 0.78 $\pm$ 0.06
MultipleModels-TriMentoring 0.86 $\pm$ 0.03 0.86 $\pm$ 0.06 0.9 $\pm$ 0.04 0.73 $\pm$ 0.02 0.78 $\pm$ 0.06
MultipleModels 1.07 $\pm$ 0.06 1.01 $\pm$ 0.03 0.84 $\pm$ 0.03 0.7 $\pm$ 0.05 1.19 $\pm$ 0.04
ParetoFlow 1.46 $\pm$ 0.03 1.19 $\pm$ 0.1 1.46 $\pm$ 0.14 1.31 $\pm$ 0.1 0.71 $\pm$ 0.05
PGD-MOO + DataPruning (Ours) 0.66 $\pm$ 0.08 0.61 $\pm$ 0.03 0.6 $\pm$ 0.03 0.6 $\pm$ 0.04 0.8 $\pm$ 0.05
PGD-MOO (Ours) 0.68 $\pm$ 0.07 0.78 $\pm$ 0.09 0.65 $\pm$ 0.03 0.6 $\pm$ 0.03 0.76 $\pm$ 0.03
Table 9: Evaluation of the $\Delta$-spread metric with 256 sampled designs on subsets of the RE task. Results are averaged over 5 different random seeds. Lower values are better.
Method RE21 RE22 RE23 RE24 RE25 RE31 RE32 RE33
MultiHead-GradNorm 0.77 $\pm$ 0.15 1.61 $\pm$ 0.37 1.7 $\pm$ 0.35 0.98 $\pm$ 0.4 1.54 $\pm$ 0.5 0.91 $\pm$ 0.17 1.26 $\pm$ 0.28 0.92 $\pm$ 0.14
MultiHead-PcGrad 0.47 $\pm$ 0.04 1.8 $\pm$ 0.22 1.14 $\pm$ 0.21 1.25 $\pm$ 0.4 1.6 $\pm$ 0.28 1.05 $\pm$ 0.25 1.09 $\pm$ 0.13 0.91 $\pm$ 0.19
MultiHead 0.42 $\pm$ 0.04 1.43 $\pm$ 0.22 1.03 $\pm$ 0.23 1.13 $\pm$ 0.19 1.86 $\pm$ 0.04 0.94 $\pm$ 0.14 0.84 $\pm$ 0.14 1.16 $\pm$ 0.05
MultipleModels-COM 0.56 $\pm$ 0.12 1.22 $\pm$ 0.42 1.2 $\pm$ 0.49 1.08 $\pm$ 0.23 1.63 $\pm$ 0.1 1.24 $\pm$ 0.17 1.23 $\pm$ 0.19 0.87 $\pm$ 0.07
MultipleModels-ICT 0.38 $\pm$ 0.02 1.78 $\pm$ 0.12 0.84 $\pm$ 0.07 0.67 $\pm$ 0.12 1.54 $\pm$ 0.04 1.28 $\pm$ 0.19 0.91 $\pm$ 0.09 1.0 $\pm$ 0.24
MultipleModels-IOM 0.98 $\pm$ 0.4 1.84 $\pm$ 0.08 1.26 $\pm$ 0.23 1.58 $\pm$ 0.24 1.54 $\pm$ 0.25 1.07 $\pm$ 0.27 1.09 $\pm$ 0.19 0.99 $\pm$ 0.06
MultipleModels-RoMA 1.19 $\pm$ 0.09 1.57 $\pm$ 0.56 1.33 $\pm$ 0.49 1.22 $\pm$ 0.24 1.44 $\pm$ 0.43 1.39 $\pm$ 0.27 1.15 $\pm$ 0.18 0.89 $\pm$ 0.06
MultipleModels-TriMentoring 0.37 $\pm$ 0.01 1.73 $\pm$ 0.12 0.91 $\pm$ 0.15 0.56 $\pm$ 0.03 1.35 $\pm$ 0.04 1.17 $\pm$ 0.14 0.99 $\pm$ 0.09 0.85 $\pm$ 0.03
MultipleModels 0.37 $\pm$ 0.02 1.85 $\pm$ 0.12 0.82 $\pm$ 0.46 0.99 $\pm$ 0.22 1.72 $\pm$ 0.36 1.65 $\pm$ 0.22 0.9 $\pm$ 0.26 1.16 $\pm$ 0.2
ParetoFlow 1.5 $\pm$ 0.12 1.37 $\pm$ 0.11 - - - 1.66 $\pm$ 0.03 1.34 $\pm$ 0.0 1.07 $\pm$ 0.11
PGD-MOO + Data Pruning (Ours) 0.61 $\pm$ 0.03 1.29 $\pm$ 0.08 1.42 $\pm$ 0.11 1.13 $\pm$ 0.01 1.12 $\pm$ 0.08 1.34 $\pm$ 0.28 1.62 $\pm$ 0.1 0.83 $\pm$ 0.17
PGD-MOO (Ours) 0.61 $\pm$ 0.03 1.28 $\pm$ 0.07 1.08 $\pm$ 0.06 1.14 $\pm$ 0.02 1.17 $\pm$ 0.07 1.32 $\pm$ 0.15 1.59 $\pm$ 0.04 0.89 $\pm$ 0.06
Table 10: Evaluation of the $\Delta$-spread metric with 256 sampled designs on subsets of the RE task. Results are averaged over 5 different random seeds. Lower values are better.
Method RE34 RE35 RE36 RE37 RE41 RE42 RE61
MultiHead-GradNorm 0.96 $\pm$ 0.23 1.17 $\pm$ 0.14 1.19 $\pm$ 0.19 1.19 $\pm$ 0.51 1.13 $\pm$ 0.5 1.12 $\pm$ 0.51 0.72 $\pm$ 0.05
MultiHead-PcGrad 1.11 $\pm$ 0.08 0.99 $\pm$ 0.13 0.91 $\pm$ 0.17 0.76 $\pm$ 0.04 0.62 $\pm$ 0.01 0.9 $\pm$ 0.08 0.72 $\pm$ 0.03
MultiHead 1.18 $\pm$ 0.02 1.03 $\pm$ 0.06 1.15 $\pm$ 0.16 0.76 $\pm$ 0.03 0.64 $\pm$ 0.02 0.86 $\pm$ 0.05 0.74 $\pm$ 0.06
MultipleModels-COM 1.16 $\pm$ 0.02 0.89 $\pm$ 0.03 0.94 $\pm$ 0.13 0.73 $\pm$ 0.02 0.56 $\pm$ 0.02 0.68 $\pm$ 0.07 0.66 $\pm$ 0.03
MultipleModels-ICT 1.06 $\pm$ 0.06 1.09 $\pm$ 0.04 1.05 $\pm$ 0.04 0.75 $\pm$ 0.04 0.61 $\pm$ 0.04 0.75 $\pm$ 0.04 0.65 $\pm$ 0.05
MultipleModels-IOM 1.09 $\pm$ 0.05 1.03 $\pm$ 0.06 1.15 $\pm$ 0.05 0.67 $\pm$ 0.04 0.57 $\pm$ 0.02 0.73 $\pm$ 0.06 0.62 $\pm$ 0.08
MultipleModels-RoMA 1.02 $\pm$ 0.03 1.22 $\pm$ 0.07 0.93 $\pm$ 0.13 0.71 $\pm$ 0.02 0.59 $\pm$ 0.01 0.66 $\pm$ 0.05 0.59 $\pm$ 0.05
MultipleModels-TriMentoring 1.05 $\pm$ 0.14 0.82 $\pm$ 0.11 1.2 $\pm$ 0.24 0.74 $\pm$ 0.02 0.58 $\pm$ 0.01 0.78 $\pm$ 0.07 0.63 $\pm$ 0.05
MultipleModels 1.22 $\pm$ 0.03 1.07 $\pm$ 0.12 1.03 $\pm$ 0.07 0.82 $\pm$ 0.03 0.62 $\pm$ 0.04 0.83 $\pm$ 0.08 0.7 $\pm$ 0.06
ParetoFlow 0.88 $\pm$ 0.05 1.13 $\pm$ 0.02 1.05 $\pm$ 0.02 1.12 $\pm$ 0.1 1.11 $\pm$ 0.07 0.9 $\pm$ 0.1 0.68 $\pm$ 0.02
PGD-MOO + Data Pruning (Ours) 0.58 $\pm$ 0.05 0.67 $\pm$ 0.04 0.7 $\pm$ 0.05 0.44 $\pm$ 0.03 0.42 $\pm$ 0.0 0.49 $\pm$ 0.03 0.52 $\pm$ 0.03
PGD-MOO (Ours) 0.56 $\pm$ 0.02 0.9 $\pm$ 0.12 0.64 $\pm$ 0.04 0.5 $\pm$ 0.01 0.43 $\pm$ 0.02 0.49 $\pm$ 0.04 0.58 $\pm$ 0.04